solution A is 2% alcohol. solution B is 6% alcohol A service station owner wants to mix the two to get 60 liters of a solution that is 3.2% alcohol How many liters of each should he mix
Answer by stanbon(75887)

solution A is 2% alcohol. solution B is 6% alcohol A service station owner wants to mix the two to get 60 liters of a solution that is 3.2% alcohol How many liters of each should he mix
-----
Equation:
alcohol + alcohol = alcohol
0.02x + 0.06(60-x) = 0.032*60
--------
Multiply thru by 1000 to get:
20x + 60*60 - 60x = 32*60
-40x = -28*60
x = 42 liters (amt. of 2% solution needed)
60-x = 18 liters (amt. of 6% solution needed)
------------------------------------------
Cheers,
Stan H.
=================